OpenStudy (anonymous):

SOLVE (2x-5/(x-2)^2=3/(x+2)

OpenStudy (anonymous):

\[(2x-5/(x-2)^2=3/(x+2)\]

OpenStudy (anonymous):

ok well i would distribute the exponent first

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

that gives you this correct? (2x−5/(x−2)(2x−5/(x−2)=3/(x+2)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now get rid of the denominator

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

Idk? That's why i'm asking for help.

OpenStudy (anonymous):

first off where does your missing parenthesis go?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):



OpenStudy (anonymous):

No. Oh so the 2x-5 is over x-2 and all of that has an exponent of two. And then that is equal to 3 over x+2

OpenStudy (anonymous):

so (2x−5/(x−2))^2=3/(x+2)

OpenStudy (tamtoan):

icey, the way you wrote down the problem is confusing :) it's missing something or something is not there :)

OpenStudy (anonymous):

sorry ((2x−5)/(x−2))^2=3/(x+2)

OpenStudy (anonymous):

Nick, that's wha tI wrote down. and i don't think anything is missing.

OpenStudy (anonymous):

that minus a couple parenthesis which makes a difference

OpenStudy (tamtoan):

is it [2x-5/(x-2)]^2=3/(x+2) or 2x-5/(x-2)^2=3/(x+2)

OpenStudy (anonymous):

unless your in a really advanced math class i think you messed something up

OpenStudy (anonymous):

No Nick you had it correct.;

OpenStudy (anonymous):

x~~-1.42875-1.11022x10^-16 i x~~2.35334+2.22045x10^-16 i x~~2.82541-2.22045x10^-16 i these are your answers so i think you messed something up

OpenStudy (anonymous):

2x-5/(x-2)^2=3/(x+2) which is the correct equation ??? : \(\large \frac{2x-5}{(x-2)^2}=\frac{3}{x+2} \) or is it: \(\large (\frac{2x-5}{x-2})^2=\frac{3}{x+2} \)

OpenStudy (anonymous):

second one

OpenStudy (tamtoan):

yeah, thank you byteme :) i wish i know how to write like you do.

OpenStudy (anonymous):

since it's the second one then: \(\large (\frac{2x-5}{x-2})^2=\frac{3}{x+2} \) \(\large \frac{(2x-5)^2}{(x-2)^2}=\frac{3}{x+2} \) \(\large (2x-5)^2(x+2)^2=3(x-2)^2 \) just "cross multiply" yuck... that left side is a 4th degree polynomial....

OpenStudy (anonymous):

my bad... actually a 3rd degree polynomial. it should be: \(\large (2x-5)^2(x+2)=3(x-2)^2 \)

OpenStudy (anonymous):

2x^2+4x-5x-10=3x^2+12+6x => x^2+7x+2 => X1=(-7V41)/2 or X2=(-7V42)/2